OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5 over 9 (F - 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 2 degrees Celsius is.

OpenStudy (anonymous):

c = 9/5 (f - 32) 5/9 c = (f - 32) 5/9 c +32 =F

OpenStudy (anonymous):

now put c = 2 in abv relation you get F = 33.33

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

@Niall?

OpenStudy (anonymous):

Gimme a sec. I think I understand. :)

OpenStudy (anonymous):

I don't think your right, cause http://www.csgnetwork.com/tempconvjava.html says that 2 c is 36 f. :)

OpenStudy (anonymous):

@kathy0514

OpenStudy (anonymous):

thats wierd

OpenStudy (anonymous):

I know

OpenStudy (cggurumanjunath):